# GuideGen: A Text-Guided Framework for Full-torso Anatomy and CT Volume Generation

### v2

@@ 4.2 Ablation studies
      Method       DSC       $\text{HD}_{95}$
      LU       CO       KI       Avg.       LU       CO       KI       Avg.
      Real       0.69       0.47       0.72       0.63       11.9       212.7       70.9       98.5
      MONAI-GM       0.48       0.10       0.24       0.27       31.7       284.9       293.5       203.4
      Zhuang’s       0.12       0.07       0.13       0.11       716.0       274.1       511.5       500.5
      MedDDPM       0.10       0.09       0.29       0.16       776.0       273.6       119.4       389.7
      MedSyn       0.44       0.11       0.39       0.31       33.1       267.4       309.0       203.2
      GuideGen       0.69       0.24       0.51       0.48       10.5       244.0       154.5       136.3
Table 6: Segmentation performance on tumor segmentation tasks of different organs using real or synthetic cases from each method.
         Mask Synthesizers          LPIPS          FID          FVD128
         None          0.56          279          5981
         LDM          0.40          61.4          2217
         MedGen3D          0.41          69.6          2045
         TCSS          0.24          20.5          785
Table 7: CT generation quality with different mask synthesizers. “None” denotes that the mask condition is removed (only the text condition is used).
GuideGen Settings generative metrics usability metrics
CAE LFG LPIPS FID FVD128 Avg. DSC Avg. HD95
N N 0.35 48.1 1729 0.29 201.9
N Y 0.36 46.4 1476 0.30 215.1
Y N 0.25 23.1 1004 0.37 180.4
Y Y 0.24 20.5 785 0.48 136.3
Table 8: Ablation studies on CAE and LFG modules. We use “Y”/“N” in CAE to indicate whether a contrast-aware rescaler is in effect and “Y”/“N” in LFG to indicate whether we use cross-attention in semantic guidance described in Sec. 3.3 as opposed to direct mask downsampling as in [8].
We evaluate the effectiveness of each component in GuideGen, including the semantic synthesizer in TCSS, contrast-aware module in CAE and the efficacy of LFG as opposed to using latent diffusion models with straightforward mask downsampling. The usability metrics are computed in the same way as Tab. 6. From Tab. 7, we observe that CT synthesized using TCSS-generated semantics performs the best, and a lack of semantic guidance degrades performance drastically. Comparing Tab. 7 with Tab. 1, we can see an interesting trend that generated image quality is roughly in line with the quality of generated masks. Tab. 8 shows a clear boost in generative metrics by integrating contrast-awareness to image autoencoders for full-torso CT generation. The LFG strategy also provides for the image quality, and more importantly it benefits in preserving tumor semantics, as shown by the increase in usability metrics. Similarly, CAE also boosts performance in the downstream usability, we suspect this comes as a collateral benefit brought by an increased quality of CT synthesis.
@@ 5 Conclusion
In this paper, we have demonstrated a novel framework for generating paired 3D anatomies and 3D CT images covering the entire torso, conditioned on a free-text medical prompt. Firstly, we introduce a text-conditional semantic synthesizer with categorical modeling to aid in the text-based synthesis of semantic masks. Secondly, we propose a contrast-aware autoencoder to accommodate the high dynamic range of CT images. Finally, we propose a latent-guided feature generator to promote detail preservation and textual alignment over conventional resampling methods. Our model outperforms a range of existing methods under comprehensive evaluation, including generative quality, alignment with textual and semantic guidance, and usability in training semantic segmentation models. Overall, our approach not only advances the state of the art in text-guided 3D CT generation but also lays a foundation for more robust training dataset synthesis for medical image analysis, making it highly impactful for clinical applications.
Supplementary Material
6 Appendix
Code Disclosure Statement
All source code, include model training, data processing, downstream evaluation and metric computation are contained in the code.zip submitted along with this appendix, and they will be publicly released should our paper be accepted.
6.1 Datasets and Implementation Details
6.1.1 Dataset Details
@@
We compiled 12 TCIA [12] datasets on three common types of cancer (lung, kidney and colon). To supplement the scarcity of publicly available colon cancer multimodal (paired CT and clinical information) dataset, we also include one private real-world dataset, dubbed RJ, for colorectal cancer. The preprocessing of all the collected unstandardized samples include the following four steps: (a) Data cleansing: All of the collected CT samples are reoriented into RAS orientation and resampled to isotopic spacing, with duplicate, impaired, partial DICOM series identified and removed. (b) Label Extraction: CT images in the collected dataset are segmented by TotalSegmentator [63] and pretrained models [34] on their respective tumor types to retrieve both multi-organ and tumor segmentation masks; their corresponding clinical information are retrieved using a private medical GPT in a certain format: The patient is {demographics group}. In this imaging, the patient’s condition is described as {clinical information}. (c) Noise Removal: Organ labels belonging to the same anatomical class (e.g. individual rib label and the anatomical class ribs) are merged into one semantic label to mitigate noise in the multi-organ labels. Samples with no or erroneous (tumor predictions outside organ prediction) tumor mask predictions are discarded. (d) Dataset Preparation: We randomly split the processed images into 4534 training cases and 1179 validation cases. Specifically, we use the CT, clinical information and masks in the training split to train GuideGen as well as baseline models, detailed description of each training dataset of choice are listed in Tab. 9, and we showcase the distribution of each demographics group in TCIA and RJ datasets in Fig. 4.
@@
       CMB-CRC [5]        Colorectal Cancer        54        237        50
       TCGA-READ [43]        Rectum Cancer        3        34        4
       TCGA-COAD [44]        Colon Cancer        25        93        21
@@
       CPTAC-CCRCC [50]        Renal Cancer        262        727        248
       TCGA-KICH [48]        Renal Cancer        15        109        26
       TCGA-KIRC [1]        Renal Cancer        267        2654        734
       TCGA-KIRP [47]        Renal Cancer        33        376        56
       CPTAC-LSCC [52]        Pulmonary Cancer        212        238        236
       CPTAC-LUAD [51]        Pulmonary Cancer        244        242        71
       CMB-LCA [6]        Pulmonary Cancer        97        696        64
       TCGA-LUAD [2]        Pulmonary Cancer        69        624        120
       TCGA-LUSC [42]        Pulmonary Cancer        37        279        92
@@
We we choose the abdominal split of Multi-Atlas Labeling Beyond the Cranial Vault (BTCV) [46] and Abdominal [37] Multi-Organ Segmentation (AMOS) datasets and evaluate the segmentation performance on some of their organ labels that overlap with our interest. Additionally, for tumor segmentation, we select two Medical Segmentation Decathlon (MSD) [3] datasets along with the Kidney Tumor Segmentation (KiTS) [26] dataset, separately for pulmonary, colorectal and renal cancer. Their details are documented in Tab. 10. We also downsamples the training cases in each inference dataset to $128\times 128\times 128$ to ensure a fair comparison with the $128\times 128\times 128$ volumes generated by GuideGen.
@@
      BTCV [46]       Abdominal Organs       24       6       Spl., Kid., Liver, Sto., Pan., Adr., Eso., Aorta, IVC, Gall., PV&SV
      AMOS [37]       Abdominal Organs       240       60       Spl., Kid., Liver, Sto., Pan., Adr., Eso., Aorta, IVC, Gall., Duo., Blad.
      MSD-Lung [3]       Pulmonary Cancer       40       10       Lung Tumor
      MSD-Colon [3]       Colon Cancer       80       20       Colon Tumor
      KiTS [26]       Renal Cancer       210       90       Kidney Tumor
@@ 6.1.2 Implementation Details
In Stage-I, we use TCSS to generate 19 major pleural, abdominal and pelvic structure***The 19 major anatomical structure of choice are: Lung, Heart, Ribs, Vertebra, Esophagus, Aorta, Inferior Vena Cava, Liver, Spleen, Stomach, Duodenum, Gallbladder, Splenic Vein and Portal Vein, Kidneys, Adrenal Glands, Pancreas, Small Bowel, Colon (with rectum) and Urinary Bladder, covering major anatomies throughout the torso. and 1 tumor semantics. The reason for not choosing more organ labels lies largely in that other anatomies suffer from relatively poor multi-organ segmentation accuracies. It should be noted that our framework should be simulate as many anatomies as possible if the quality of training data is ensured. The model channels in each layer of our backbone U-Net network are 32, 64, 128, 256, respectively, with 8$\times$ downsampling at the U-Net bottleneck. Each layer consists of two ResNet convolution blocks and we add a cross-attention layer to the first and third layer to integrate textual information.
In Stage-II, our CAE downsamples the CT image by 4$\times$ and operates on a codebook of 8192 entries, each with a feature dimension of 8. The frame discriminator $\mathbf{D}_{f}$ in CAE chooses 8 slices for discriminative loss computation in each iteration. We also set a min-max normalization for CT images prior to their input into CAE model: CT images are linearly rescaled in such a way that the CT intensities of the range [-1500, 1500]HU resides in the interval [-1, 1] after rescaling. In this way, we can recover CT intensities from the decoder half of CAE after LFG generation by multiplying the result by the multiplier 1500.
In Stage-III, we first encode tumor and organ semantics separately to a vector with a feature dimension of 4 and spatial dimension 4$\times$ smaller than original input (which computes to $32\times 32\times 32$ in our experiment setting). As we jointly generate tumor and organ semantics in the first stage, we follow the following steps to retrieve label for them individually: Firstly, for 19 organ labels $l_{A}=\{l_{A}^{1},\cdots,l_{A}^{19}\}$, 1 tumor label $l_{T}$ and 1 background label $l_{B}$, the final step of TCSS sampling trajectory will output a probability vector $\hat{\mathbf{x}}_{0}\in[0,1]^{21\times H\times W\times D}$. $H$, $W$ and $D$ are spatial dimensions, which all equals 128 in our setting. Secondly, we perform argmax operation on the channel dimension to retrieve these semantic masks, specifically, the organ semantics $\mathbf{m}_{A}$ and tumor semantics $\mathbf{m}_{T}$ are retrieved by
@@
That is, we discard the probabilities of tumor label when retrieving organ semantics, and $\delta(x,y)$ here is the impulse function that takes 1 on the entry $i$ when $x[i]$ equals $y$ and 0 otherwise. The encoded semantic latents, along with the image feature encoded by the encoder half of CAE are then concatenated to a latent vector with a feature dimension of 16, which is fed into a backbone U-Net with the same structure as that described in TCSS. Loss is computed between the U-Net output and the image latent.
@@
We follow official implementation available online for each baseline method, and adjust their model parameters (mainly the channel numbers for each convolutional layer) so that they can generate $128\times 128\times 128$ volumes under our physical setting (NVIDIA RTX 4090 GPUs). We also skip the final super-resolution module in some baselines [64, 23] since it has no relevance to conditional generation and can be added to any generative framework as a postprocessing module.
6.2 Supplementary Results
@@
This section provides tables and figures that serves as a supplement to the main results displayed in the main article.
6.2.1 Qualitative Results
@@
In this subsection, we first showcase GuideGen’s generative performance of both unhealthy and healthy samples using randomly chosen demographic and clinical descriptions from TCIA and RJ datasets in Fig. 5. Secondly, we display multi-organ segmentation results and tumor segmentation results separately in Fig. 6 and Fig. 7. For slices located in the pleural region, we display them using a normalization window at [-975, -225]HU, while for slices in the abdominal and pelvic regions, we display them using a normalization window at [-50, 150]HU as recommended in [58].
6.2.2 Quantitative Results
@@
We display standard variances of each metric used in multi-organ segmentation and tumor segmentation evaluations separately in Tab. 6.2.3 and Tab. 12 in this subsection. We also report the validation accuracy and AUROC of demographic or clinical classifiers used to assess image-prompt alignment in Tab. 13. They are implemented based on the encoder half of U-Net 3D networks [11] and trained on the training split of TCIA and RJ datasets. Classifier classes are chosen following the demographic classes in Fig. 4 and tumor locations in Tab. 9. Metrics in Tab. 13 are computed on the validation split of TCIA and RJ datasets.
6.2.3 Utilizing GuideGen for Dataset Synthesis or Augmentation
Finally, in Tab. 14, we tested the usability of generated samples by using different number of GuideGen-generated cases for dataset synthesis (segmentation models trained with no real cases required) or dataset augmentation (using GuideGen-generated cases as a complement to the real cases to train segmentation models). We randomize the demographic features and clinical severity in GuideGen’s input while fixing the tumor location to a certain organ. It can be seen that on the multi-organ segmentation tasks, using 1000 synthetic cases as augmentation achieves an average performance gain of 6% across BTCV and AMOS datasets, while segmentation models trained using synthetic data alone achieves comparable performance to those trained with real data. For tumor segmentation tasks, performance gains are also observed using synthetic data as augmentation at a large scale, achieving a surplus in DSC of 9%, 1% and 5% separately for three different tumor types at 1000 synthetic cases. Using synthetic data for dataset synthesis also provides a solid starting point, especially for lung tumors, as seen in the last 6 rows of Tab. 14.

### v3

@@ 4.2 Ablation studies
[…]
Cases LU CO KI Avg. LU CO KI Avg.
Real 50/100/313 0.69 0.47 0.72 0.63 11.9 212.7 70.9 98.5
MAISI 200 0.48 0.10 0.24 0.27 31.7 284.9 293.5 203.4
Zhuang’s 200 0.12 0.07 0.13 0.11 716.0 274.1 511.5 500.5
MedDDPM 200 0.10 0.09 0.29 0.16 776.0 273.6 119.4 389.7
MedSyn 200 0.44 0.11 0.39 0.31 33.1 267.4 309.0 203.2
GuideGen 200 0.71 0.21 0.64 0.52 8.4 227.0 84.5 106.6
Table 6: Segmentation performance on 3 tumor segmentation tasks using real or synthetic cases from each method.
[…]
   w.o. mask input    0.42    54.3    -    0.32
   LDM-generated mask input    0.34    38.4    0.34    0.40
   MedGen3D-generated mask input    0.36    39.0    0.23    0.41
   w.o. knowledge injection    0.26    21.7    0.25    0.57
   w.o. anatomy preservation    0.27    32.4    0.40    0.61
   w.o. HDR accommodation    0.33    40.9    0.36    0.64
   GuideGen    0.25    20.2    0.52    0.69
Table 7: Ablations studies on GuideGen. DSC and Acc. are separately evaluated and averaged on 3 downstream tumor segmentation tasks and 4 image-prompt alignment tasks.
Figure 4: Segmentation performance using different number (0, 100, 200, 500, 1K) of GuideGen-generated samples as augmentation. Darker and lighter solid lines separately denote segmentation model trained with or without real data.
(1) Mask Synthesizers: By replacing the mask synthesizer (Stage-I) of our GuideGen with other off-the-shelf mask generators, we can ablate on the CT generative performance using different mask synthesizers, as illustrated by the first 3 rows and the last row of Tab. 7. It can be seen that CTs generated with semantic guidance from TCSS performs best, and a lack of semantic guidance degrades performance drastically. Comparing Tab. 7 with Tab. 1, we can see an interesting trend that generated CT quality is roughly in line with the quality of mask inputs. (2) Knowledge Injection: The superiority of knowledge injection as opposed to conventional cross-attention guidance is illustrated in rows 4, 7 of Tab. 7, with a significant boost of alignment and downstream segmentation metrics indicating that knowledge injection better preserves textual features. (3) Anatomy-aware HDR Autoencoder: The last 3 rows in Tab. 7 also shows a clear boost in generative metrics by integrating HDR accommodation and anatomy preservation to CT autoencoders for full-torso generation, which also promotes downstream usability as the image quality becomes more desirable. (4) Generation quantity: From Fig. 4, we can clearly see a trend that with a larger quantity of generated samples, downstream segmentation performance can be better boosted across all tasks of choice. Also, for tasks with a simpler background (like LU), the performance of segmentation models trained using only GuideGen-generated samples can be comparable to or even better than those trained with real cases only.
@@ 5 Conclusion
In this paper, we have demonstrated a novel framework for generating paired 3D anatomies and CT images covering the entire torso, conditioned on a structured medical prompt. Specifically, our GuideGen outperforms a range of existing methods under comprehensive evaluation, including generative quality, alignment with textual and semantic guidance, and usability in training semantic segmentation models. Overall, our approach not only advances the state of the art in text-guided 3D CT generation but also lays a foundation for more robust training dataset synthesis and augmentation for medical image analysis. Limitations While GuideGen achieves strong performance to synthesize full-torso semantics and CTs from structured prompts given by Large Language Models, its cannot directly generate plausible images from free-text inputs. We consider this as a potential future direction for enabling more versatile control.
6 Acknowledgements
This work was supported by National Natural Science Foundation of China (NSFC 62301311).
Appendix A Datasets and Implementation Details
A.1 Dataset Details
@@
We compiled 12 TCIA (clark2013cancer) datasets on three common types of cancer (lung, kidney and colon). To supplement the scarcity of publicly available colon cancer multimodal (paired CT and clinical information) dataset, we also include one private real-world dataset, dubbed RJ, for colorectal cancer. The preprocessing of all the collected unstandardized samples include the following four steps: (a) Data cleansing: All of the collected CT samples are reoriented into RAS orientation and resampled to isotopic spacing, with duplicate, impaired, partial DICOM series identified and removed. (b) Label Extraction: CT images in the collected dataset are segmented by TotalSegmentator (wasserthal2023totalsegmentator) and pretrained models (isensee2021nnu) on their respective tumor types to retrieve both multi-organ and tumor segmentation masks; their corresponding clinical information are retrieved using a private medical GPT in a certain format: The patient is {demographics group}. In this imaging, the patient’s condition is described as {clinical information}. (c) Noise Removal: Organ labels belonging to the same anatomical class (e.g. individual rib label and the anatomical class ribs) are merged into one semantic label to mitigate noise in the multi-organ labels. Samples with no or erroneous (tumor predictions outside organ prediction) tumor mask predictions are discarded. (d) Dataset Preparation: We randomly split the processed images into 4534 training cases and 1179 validation cases. Specifically, we use the CT, clinical information and masks in the training split to train GuideGen as well as baseline models, detailed description of each training dataset of choice are listed in Tab. 8, and we showcase the distribution of each demographics group in TCIA and RJ datasets in Fig. 5.
@@
   CMB-CRC (Cancer_Moonshot_Biobank2022)    Colorectal Cancer    54    237    50
   TCGA-READ (Kirk2016)    Rectum Cancer    3    34    4
   TCGA-COAD (Kirk2016-ty)    Colon Cancer    25    93    21
@@
   CPTAC-CCRCC (ccrcc-zv)    Renal Cancer    262    727    248
   TCGA-KICH (Linehan2016-qd)    Renal Cancer    15    109    26
   TCGA-KIRC (Akin2016-uc)    Renal Cancer    267    2654    734
   TCGA-KIRP (Linehan2016-uj)    Renal Cancer    33    376    56
   CPTAC-LSCC (cptac-hu)    Pulmonary Cancer    212    238    236
   CPTAC-LUAD (cptac-fz)    Pulmonary Cancer    244    242    71
   CMB-LCA (Cancer_Moonshot_Biobank2022-ra)    Pulmonary Cancer    97    696    64
   TCGA-LUAD (Albertina2016-nm)    Pulmonary Cancer    69    624    120
   TCGA-LUSC (Kirk2016-wa)    Pulmonary Cancer    37    279    92
@@
We we choose the abdominal split of Multi-Atlas Labeling Beyond the Cranial Vault (BTCV) (landman2015miccai) and Abdominal (ji2022amos) Multi-Organ Segmentation (AMOS) datasets and evaluate the segmentation performance on some of their organ labels that overlap with our interest. Additionally, for tumor segmentation, we select two Medical Segmentation Decathlon (MSD) (antonelli2022medical) datasets along with the Kidney Tumor Segmentation (KiTS) (heller2023kits21) dataset, separately for pulmonary, colorectal and renal cancer. Their details are documented in Tab. 9. We also downsamples the training cases in each inference dataset to $128^{3}$ to ensure a fair comparison with the $128^{3}$ volumes generated by GuideGen and other baselines.
@@
   BTCV (landman2015miccai)    Abdominal Organs    24    6    Spl., Kid., Liver, Sto., Pan., Adr., Eso., Aorta, IVC, Gall., PV&SV
   AMOS (ji2022amos)    Abdominal Organs    240    60    Spl., Kid., Liver, Sto., Pan., Adr., Eso., Aorta, IVC, Gall., Duo., Blad.
   MSD-Lung (antonelli2022medical)    Pulmonary Cancer    40    10    Lung Tumor
   MSD-Colon (antonelli2022medical)    Colon Cancer    80    20    Colon Tumor
   KiTS (heller2023kits21)    Renal Cancer    210    90    Kidney Tumor
@@ A.2 Implementation Details
[…]
In Stage-III, we first retrieve the semantic information for organ and tumor masks individually: Firstly, for 19 organ labels $l_{A}=\{l_{A}^{1},\cdots,l_{A}^{19}\}$, 1 tumor label $l_{T}$ and 1 background label $l_{B}$, the final step of TCSS sampling trajectory will output a probability vector $\hat{\mathbf{x}}_{0}\in[0,1]^{21\times H\times W\times D}$. $H$, $W$ and $D$ are spatial dimensions, which all equals 128 in our setting. Secondly, we perform argmax operation on the channel dimension to retrieve these semantic masks. Specifically, we find that it is better to retrieve the semantic labels via the following way separately for organ $\mathbf{m}_{A}$ and tumor semantics $\mathbf{m}_{T}$:
@@
That is, we discard the probabilities of tumor label when retrieving organ semantics. This is done so that the diffusion model can accurately pin each tumor type on a specific organ and to raise special awareness to our latent-guided generator by separating tumor mask as a separate semantic channel. $\delta(x,y)$ here is the impulse function that takes 1 on the entry $i$ when $x[i]$ equals $y$ and 0 otherwise. These semantic latents ($\mathbf{m}_{A}$ and $\mathbf{m}_{T}$) are then resampled to a desired output resolution (in our case it is 64 following the latent spatial resolution of our anatomy-aware HDR autoencoder), along with the image feature encoded by the encoder half of our autoencoder are then concatenated to a latent vector with a feature dimension of 16, which is fed into a backbone U-Net with the same structure and knowledge injection module as that described in TCSS. Loss is computed between the U-Net output and the noised image latent.
@@
We follow the official implementations available online for each baseline method, and adjust their model parameters (mainly the channel numbers for each convolutional layer) so that they can generate $128^{3}$ volumes under our physical setting (NVIDIA RTX 4090 GPUs). We did not use baseline models to generate $256^{3}$ volumes here because it would compromise some models’ performance drastically and instead we adapt GuideGen to generate $128^{3}$ volumes to compare with them fairly (see the Experiment section in the main text). We also skip the final super-resolution stage in some baselines (xu2024medsyn; hamamci2025generatect) since it has no relevance to conditional generation and can be added to any generative framework as a postprocessing module.
Appendix B Supplementary Results
@@
This section provides tables and figures that serve as a supplement to the main results displayed in the main article.
B.1 Qualitative Results
@@
In this section, we first showcase GuideGen’s generative performance of both unhealthy and healthy samples using randomly chosen demographic and clinical descriptions from TCIA and RJ datasets in Fig. 6. Secondly, we display multi-organ segmentation results and tumor segmentation results separately in Fig. 7 and Fig. 8. For slices located in the pleural region, we display them using a normalization window at [-975, -225]HU, while for slices in the abdominal and pelvic regions, we display them using a normalization window at [-50, 150]HU as recommended in (seeram2015computed). We generate mask volumes $\hat{\textbf{m}}$ at a resolution of $128^{3}$ and use trilinear interpolation to resample them to $64^{3}$ for CT volume generation. The generated $256^{3}$ CT volume $\hat{\textbf{v}}$ is then paired with $\hat{\textbf{m}}$ upsampled to the same spatial resolution as $\hat{\textbf{v}}$ as generated pair.
B.2 Quantitative Results
@@
We display standard variances of each metric used in multi-organ segmentation and tumor segmentation evaluations separately in Tab. C and Tab. 11 in this section. We also report the validation accuracy and AUROC of demographic or clinical classifiers used to assess image-prompt alignment in Tab. 12. They are implemented based on the encoder half of U-Net 3D networks (cciccek20163d) and trained on the training split of TCIA and RJ datasets. Classifier classes are chosen following the demographic classes in Fig. 5 and tumor locations in Tab. 8. Metrics in Tab. 12 are computed on the validation split of TCIA and RJ datasets.
B.3 Comparison with Tumor Inpainting Methods for Segmentation Augmentations
Finally, it has come to our knowledge that in addition to full-image generation like GuideGen and baseline methods described in the main paper, there has been a growing interest in using image inpainting to synthesize only lesion or tumor sites, such as DiffTumor (chen2024towards), FreeTumor (wu2024freetumor) and LesionDiffusion (tian2025lesiondiffusion). Here, we hope to first emphasize that GuideGen, enabled by TCSS, can synthesize not only tumor masks, but also full anatomies and that it learns the tumor distribution through learning, rather than random deformation as per current practices. Nevertheless, to illustrate that GuideGen is also capable of synthesizing high-quality tumors, we perform segmentation pretraining using 1) real CTs and inpainted tumor patches via various tumor generation pipelines and 2) GuideGen-generated whole CT volumes and present the results in Tab. 13. It is evident that GuideGen can synthesize faithful tumor sites that aid for downstream tasks and are comparable or even superior to two baseline tumor inpainting methods’ synthesized cases, even if these inpainting methods comes with task-specific designs. This comparison further highlights GuideGen’s competence as a segmentation dataset generator.
Appendix C Limitation and Future Directions
Although GuideGen currently can achieve downstream segmentation dataset construction and augmentation, its generation pipeline contains three-stages, which can make adaptation to other modalities or out-of-distribution datasets time-consuming. An end-to-end model would certainly be more preferable, but due to the ambiguity problem mentioned in the Method section in the main text, using a single diffusion model with continuous modeling may be insufficient to fully grasp the exact discrete nature of labels. Future works may revolve around finding more efficient way to represent discrete labels, which may include using codebooks to encode adjacent labels into distant vectors in high-dimensional space, thereby circumventing the issue of ambiguity. Another limitation of our work is that the input prompt, despite having no constraint on the words of choice, needs to be input with a certain format to draw the model’s attention. A possible future direction can be using vision-language foundation models (carvalhido2025stress) to align features extracted from structure-free inputs with encoded image latents.